OpenStudy (anonymous):

what is the solution in the set of real numbers of the quadratic equation \[x ^{2}-3x+10=0\]

OpenStudy (dumbcow):

use quadratic formula a=1 b=-3 c=10 \[\frac{-b \pm \sqrt{b ^{2}-4ac}}{2a}\] this gives 3 +-sqrt(-31)/2 =3/2 +- sqrt(31)/2 i